OpenStudy (anonymous):

Help complete the square f(x)=-x^2-4x-8

OpenStudy (anonymous):

start with \(f(x)=-(x^2+2x)-18\) then go to \[f(x)=-(x+1)^2+k\] to find \(k\) replace \(x\) in the original function by \(-1\) i.e. \(k=f(-1)\)

OpenStudy (anonymous):

\(f(-1)=?\)

OpenStudy (anonymous):

to me it looks like it would be -1

OpenStudy (anonymous):

oops \[f(-1)=-(-1)^2-2\times (-1)-18=-1+2-18=-17\] so \[f(x)=-(x+1)^2-17\]

OpenStudy (anonymous):

f(x)=-(x-1)^2-17

OpenStudy (anonymous):

should be plus one, not minus one

OpenStudy (anonymous):

\[f(x)=-(x+1)^2-17\] you can check by multiplying out to see that it is right